How to find the value of sin4x,cos4x,cot4x,or more 3x . Please its an urgent?
 

Grade:11

3 Answers

Vijay Mukati
8 years ago
Dear Student, If you know the value of sinx or cosx or tanx etc. Then you can easly find the values of sin4x, cos4x, cot4x by using the basic trignometric formulas.
Example:- sin 4x = sin2(2x) = 2sin2xcos2x=4sinxcosxcos2x. Also you can use cos2x = 2cos2x-1 or 1-2sin2x formula. Thanks.
VINEETH N
8 years ago
What will be the final result of sin4x? Is it 2cos2x? How?
 
Vijay Mukati
8 years ago
sin 4x = sin2(2x) = 2sin2xcos2x=4sinxcosxcos2x = 4sinxcos(2cos2x-1) =4sinxcos(1-2sin2x).